# Software, Formal analysis, Data Curation

### 24 Dec 2023 v2

@@ Abstract
Methods: A systematic examination of 152,735 repositories yielded 2,010 relevant hits. We then analyzed 538 relevant commits and 208 relevant issues using inductive and deductive coding and corroborated findings with discussions from Stack Overflow.
@@ 1 Introduction
  • we triangulate this information extraction by identifying Stack Overflow posts where pertinent discussions take place;
@@ 3.3.1 RQ1
To answer the first research question, we must identify what kind of information related to cloud cost management can be extracted from the commits. Out of this set of 6,116 commits, $377$ are from forked repositories already in the set. After filtering out these commits and those that do not modify any Terraform files, we are left with 1,162 repositories and 2,045 related commits. The selected commits are then inspected manually to decide their actual relevance. For that, each commit message is checked by two researchers and validated by a third one. Any conflicts are resolved by the entire team in consolidation meetings. The output of this process results in the identification of $538$ relevant commits. The selected units come from $434$ distinct repositories.
In Figure 2, we characterize the repositories in terms of IaC artifacts’ size (iac-sloc, max. $89$K), source code size (sloc, max. 1 21012101\,2101 210K), number of commits (#commits, max. $17$K) and contributors (#contributors, max. $596$). The max values are not visible in Figure 2 since we needed to trim the top $10\%$ values of each variable to better visualize the plot. We also note that $65\%$ ($284$) of the population comprise repositories that contain (and manage) IaC artifacts only (i.e., they do not contain source code). These characteristics suggest that the population is varied with a tendency for repositories maintained by few contributors and concentrated in repositories of more Terraform and other IaC files than source code. This is possibly an indication of developers taking over a “Terraform expert” role in teams that become responsible also with dealing with the operational cost aspect of each project.
@@ 3.3.2 RQ2
For the second research question, we wish to explore what kind of additional information related to cloud cost awareness can be extracted from entries in issue trackers that adds to that extracted from commit messages. As shown in Figure 1, we start from the list of 1,339 repositories obtained from the first filtering of commits based on keyword. We use this list because discussion may take place before actual changes happen (as demonstrated by modification to Terraform files), and we would like to capture them too. Also, this set of repositories allows us to narrow down the population to an amount of issues that we can feasibly extract from GitHub since they already contain some evidence that cost may be a concern for the project.
@@
Next, we apply a process similar to that described for RQ1 to filter for relevant issues and then label them. Since issues may contain long discussions, the coding is focused on the context, i.e., the sentence where the keyword appears and the surrounding ones (when needed). If multiple keywords appear on the same issue and they refer to different contexts, multiple checks are performed. We note that the labels are applied to whole issues based on evidence found in any of the content elements (i.e., title, body or comments). This process culminates in the selection of $208$ units belonging to $89$ distinct repositories. In Figure 3, we characterize the repositories in terms of IaC size (iac-sloc, max. $43$K), source code size (sloc, max. 1 48614861\,4861 486K), and number of commits (#commits, max. $6$K) and contributors (#contributors, max. $156$). We again trimmed the top $10\%$ values of each variable to better visualize the plot. This population of repositories is considerably smaller than that collected for RQ1, and contains a lower amount of projects that contain (and manage) Terraform files (only $44\%$). Nevertheless, the descriptive statistics summarized by Figure 3 suggest a fair distribution of data points, with higher averages (by approx. a factor of two).
@@ 4.1 Cost Awareness in Commits (RQ1)
In this respect, this label can be located anywhere between the previous two ones: it can identify awareness and intention of action at the same time. On a related note, it is worth point out that, although less prevalent, we also identified information explicitly discussing cost increases due to a prior change (in roughly $2\%$ of the units).
@@ 4.3 Knowledge Organization (RQ3)
In particular, only one topic derived from issues was considered. The relevant topics mention cost-related terms (e.g., ‘cheap’, ‘expensive’, ‘budget’, ‘waste’) and actions (e.g., ‘change’, ‘move’, ‘add’, ‘test’, ‘upgrade’) associated with various properties of the deployment, both general (e.g., ‘VM’, ‘storage’, ‘disk’, ‘machine’) and specific (e.g., ‘dynamodb’, ‘CPU’, ‘NAT’, ‘EC2’, ‘RAM’). Some of the connections revealed through the topics are already observable in our coding for RQ1 and RQ2, in the form of co-occurring labels as summarized by Figure 7. The figure presents these co-occurence relationships as an UpSet plot Lex et al. (2014). The labels (with their frequencies) are shown as rows on the bottom part, and the frequency of the various combinations of labels are represented as columns on the top part and described through the connected dots on the bottom part.
@@ 5.1 Data Collection
The data collection entailed the extraction of relevant discussions from SO. For that, we used key datasets from the September 12, 2023 data dump, namely, posts [^12] https://archive.org/download/stack-exchange-data-dump-2023-09-12/stackoverflow.com-Posts.7z, comments [^13] https://archive.org/download/stack-exchange-data-dump-2023-09-12/stackoverflow.com-Comments.7z and histories [^14] https://archive.org/download/stack-exchange-data-dump-2023-09-12/stackoverflow.com-PostHistory.7z, totaling 57.3 GiB. We specifically targeted discussions about Terraform, rooted in questions containing tags with the string ‘terraform’ in the post’s metadata. Thus, we started by filtering questions fitting the mentioned criteria, which led to a dataset of 19,139 questions stored as JSON files.
We gathered the comments and post histories associated with each question and updated the respective JSON file to include them. We then extracted the answers linked to each question, identified through their parent post IDs, and added the comments and post histories similarly to the questions. The answers (and associated data) were then added to the question files, creating the final version of the dataset of 19,139 questions and their associated answers, complete with their respective comments and post histories.
@@ 5.3 Results and Discussion
The initial filtering (using keywords stems) returned a total of 491 questions, i.e., approx. 2.6% of all questions on SO with ‘terraform’ in any of the tags. Although the sample size may seem small compared to the entire population of Terraform-related questions, we note that this is double the sample size compared to that of GitHub repositories containing cost-related commits (1.3%, 2,010 out of 152,735 containing Terraform artifacts; see Section 3.3). This proportional increase may be further indicative of developers’ involvement with and attention to cost-related decisions.
Upon inspecting these questions and their answers for properties from our knowledge graph, we learned that 451 (approx. 92%) of them mention one or more of these properties. In Figure 9, we show the exact number of questions where each property was identified. From these questions’ JSON file, we then extracted 3,934 sentences containing one or more of the properties. Aided by these highlighted sentences, we proceeded to manually inspect a representative sample of 37151515https://www.calculator.net/sample-size-calculator.html?type=1&cl=95&ci=5&pp=2.6&ps=491&x=Calculate questions, considering that we selected 2.6% of a population of 19,139 questions with Terraform tags.
During the inspection, we aimed to verify whether: (1) the question indeed contained a cost-related discussion (either as the main problem or as the element of a description or argument), and (2) one of the properties is an essential element of the identified cost-related text. As a result, we found that all inspected questions were true positives. These combined observations show that the knowledge graph cannot only be used to identify actual pain points faced by cloud application developers, but it also points to the most recurrent ones. Looking at the questions for which the properties do not appear, there is an opportunity to further augment the knowledge graph with missing properties, but that is left as future work.
@@
Figure 9: Number of Stack Overflow questions and their associated answers mentioning properties (low-level elements) of the knowledge graph.
@@ 6 Implications & Future Work
  • Use the extracted posts from Stack Overflow to extend the list of topics related to cost awareness, and update the knowledge graph of Fig. 8 accordingly.
@@ 7 Threats to Validity
External Validity: Regarding external validity, the population in our dataset may not represent all possible cost-related discussions in cloud projects. The inclusion of repositories hosted in other (closed-source) platforms, or that use other cloud orchestrators, or that use no cloud orchestrators, could lead to the identification of new discussions. However, our decisions were carefully considered with the aim of ensuring data quality, diversity and quantity while keeping the execution feasible for the available human resources. Furthermore, our goal was not to collect all possible evidence, to begin with, but rather to see if there is any evidence available.
Construct Validity: The selection criteria, and most noticeably the defined keywords may threat the construct validity of our work. We mitigated this threat by piloting and testing our selection criteria, and considering the knowledge of both academic and industrial domain experts. Furthermore, the coding activity is naturally open to subjectivity and inconsistencies. To mitigate this threat, we followed a well-established process, and introduced an extra final step to consolidate the knowledge and systematically discuss the labels. Finally, the topics derived from applying LDA may not fully represent their content. We mitigate this threat by manually inspecting the units of analysis during the investigation of RQ3, which was part of the axial coding and selective coding.

### 11 Apr 2024 v3

@@ Abstract
Methods: A systematic examination of 152 735152735152\,735152 735 repositories yielded 2 01020102\,0102 010 relevant hits. We then analyzed 538 relevant commits and 208 relevant issues using inductive and deductive coding and corroborated findings with discussions from Stack Overflow.
@@ 1 Introduction
  • we triangulate and augment this information extraction by identifying Stack Overflow posts where pertinent discussions take place;
@@ 3.3.1 RQ1
To answer the first research question, we must identify what kind of information related to cloud cost management can be extracted from the commits. Out of this set of 6 11661166\,1166 116 commits, $377$ are from forked repositories already in the set. After filtering out these commits and those that do not modify any Terraform files, we are left with 1 16211621\,1621 162 repositories and 2 04520452\,0452 045 related commits. The selected commits are then inspected manually to decide their actual relevance. For that, each commit message is checked by two researchers and validated by a third one. Any conflicts are resolved by the entire team in consolidation meetings. The output of this process results in the identification of $538$ relevant commits. The selected units come from $434$ distinct repositories.
@@
In Figure 2, we characterize the repositories in terms of IaC artifacts’ size (iac-sloc, max. $89$K), source code size (sloc, max. 1 21012101\,2101 210K), number of commits (#commits, max. $17$K), contributors (#contributors, max. $596$), number of Terraform modules (#tfmodules, max. $1.4$K), resources (#tfresources, max. $2.5$K), variables (#tfvariables, max. $5$K), subnets-related definitions (#tfsubnets, max. $32$), instances-related definitions (#tfinstances, max. $36$), and number of Terraform clusters-related definitions (#tfclusters, max. $21$). The max values are not visible in Figure 2 since we needed to trim the top $10\%$ values of each variable to better visualize the plot. We also note that $65\%$ ($284$) of the population comprise repositories that contain (and manage) IaC artifacts only (i.e., they do not contain source code). These characteristics suggest that the population is varied with a tendency for repositories maintained by few contributors and concentrated in repositories of more Terraform and other IaC files than source code. This is possibly an indication of developers taking over a “Terraform expert” role in teams that become responsible also with dealing with the operational cost aspect of each project.
@@ 3.3.2 RQ2
For the second research question, we wish to explore what kind of additional information related to cloud cost awareness can be extracted from entries in issue trackers that adds to that extracted from commit messages. As shown in Figure 1, we start from the list of 1 33913391\,3391 339 repositories obtained from the first filtering of commits based on keyword. We use this list because discussion may take place before actual changes happen (as demonstrated by modification to Terraform files), and we would like to capture them too. Also, this set of repositories allows us to narrow down the population to an amount of issues that we can feasibly extract from GitHub since they already contain some evidence that cost may be a concern for the project.
@@
Next, we apply a process similar to that described for RQ1 to filter for relevant issues and then label them. Since issues may contain long discussions, the coding is focused on the context, i.e., the sentence where the keyword appears and the surrounding ones (when needed). If multiple keywords appear on the same issue and they refer to different contexts, multiple checks are performed. We note that the labels are applied to whole issues based on evidence found in any of the content elements (i.e., title, body or comments). This process culminates in the selection of $208$ units belonging to $89$ distinct repositories. In Figure 3, we characterize the repositories in terms of the same metrics as in Figure 2: iac-sloc (max. $43$K), sloc (max. 1 48614861\,4861 486K), #commits (max. $6$K), #contributors, (max. $156$), #tfmodules (max. $373$K), #tfresources (max. $646$K), #tfvariables (max. $3.5$K), #tfsubnets, (max. $39$), #tfinstances (max. $34$), and #tfclusters, max. $27$). We again trimmed the top $10\%$ values of each variable to better visualize the plot. This population of repositories is considerably smaller than that collected for RQ1, and contains a lower amount of projects that contain (and manage) Terraform files (only $44\%$). Nevertheless, the descriptive statistics summarized by Figure 3 suggest a fair distribution of data points, with higher averages (by approx. a factor of two).
Figure 3: RQ2 repository demographics (90th percentile).
@@ 4.1 Cost Awareness in Commits (RQ1)
In this respect, this label can be located anywhere between the previous two ones: it can identify awareness and intention of action at the same time. On a related note, it is worth pointing out that, although less prevalent, we also identified information explicitly discussing cost increases due to a prior change (in roughly $2\%$ of the units).
@@ 4.3 Knowledge Organization (RQ3)
In particular, only one topic derived from issues was considered. The relevant topics mention cost-related terms (e.g., ‘cheap’, ‘expensive’, ‘budget’, ‘waste’) and actions (e.g., ‘change’, ‘move’, ‘add’, ‘test’, ‘upgrade’) associated with various properties of the deployment, both general (e.g., ‘VM’, ‘storage’, ‘disk’, ‘machine’) and specific (e.g., ‘dynamodb’, ‘CPU’, ‘NAT’, ‘EC2’, ‘RAM’). Some of the connections revealed through the topics are already observable in our coding for RQ1 and RQ2, in the form of co-occurring labels as summarized by Figure 7. The figure presents these co-occurence relationships as an UpSet plot (Lex et al., 2014). The labels (with their frequencies) are shown as rows on the bottom part, and the frequency of the various combinations of labels are represented as columns on the top part and described through the connected dots on the bottom part.
@@ 5.1 Data Collection
The data collection entailed the extraction of relevant discussions from SO. For that, we used key datasets from the September 12, 2023 data dump, namely, posts [^12] https://archive.org/download/stack-exchange-data-dump-2023-09-12/stackoverflow.com-Posts.7z, comments [^13] https://archive.org/download/stack-exchange-data-dump-2023-09-12/stackoverflow.com-Comments.7z and change histories [^14] https://archive.org/download/stack-exchange-data-dump-2023-09-12/stackoverflow.com-PostHistory.7z, totaling 57.3 GiB. As the goal of this process is to triangulate the results of the study on GitHub reponsitories, we specifically targeted discussions about Terraform, rooted in questions containing tags with the string ‘terraform’ in the post’s metadata. Thus, we started by filtering questions fitting the mentioned criteria, which led to a dataset of 19 1391913919\,13919 139 questions stored as JSON files.
We gathered the comments and post histories associated with each question and updated the respective JSON file to include them. We then extracted the answers linked to each question, identified through their parent post IDs, and added the comments and post histories similarly to the questions. The answers (and associated data) were then added to the question files, creating the final version of the dataset of 19 1391913919\,13919 139 questions and their associated answers, complete with their respective comments and post histories.
@@ 5.3 Results and Discussion
The initial filtering (using keywords stems) returned a total of 491 questions, i.e., approx. 2.6% of all questions on SO with ‘terraform’ in any of the tags. Although the sample size may seem small compared to the entire population of Terraform-related questions, we note that this is double the sample size compared to that of GitHub repositories containing cost-related commits (1.3%, 2 01020102\,0102 010 out of 152 735152735152\,735152 735 containing Terraform artifacts; see Section 3.3). This proportional increase may be further indicative of developers’ involvement with and attention to cost-related decisions.
@@
Figure 9: Co-occurrences of action and property concepts from the knowledge graph of Figure 8 in Stack Overflow questions and their associated answers.
Upon inspecting these questions and their answers for actions and properties from our knowledge graph, we learned that 451 (approx. 92%) of them mention one or more of these concepts. In the diagonal of Figure 9, we show the exact number of questions and associated answers where each concept was present as a term. While properties such as area appear only once, other properties such as provider and instance (189 and 138 occurrences, respectively) but also actions such as change (199) and test (154) are very popular. The same question and its associated answers might contain a reference to more than one such concept. From these questions’ JSON file, we then extracted 3 93439343\,9343 934 sentences containing one or more of these concepts.
Extracted sentences showcase developers sharing their experiences in dealing with insidious cost-inducing technicalities in using existing services or asking for advice in this direction:
Figure 10: Knowledge graph updated based on the co-occurrences of concepts in the Stack Overflow posts.
Azure by itself discusses the challenge you are addressing and recommends using ""data lifecycle"" for reducing the storage costs. [...] After you enable blob versioning for a storage account, every write operation to a blob in that account results in the creation of a new version. post id: #72881767 label: storage
I am trying to reduce the cost of my AWS infrastructure deployed using Terraform for a Django app. I have 2 public subnets and 2 private subnets and in the subnets I deploy NAT gateways and elastic ips. all works but is expensive. post id: #74422443 label: nat
Developers even identify cost-related features requested to be added to Terraform as an answer to existing problems:
Also savings plans have largely replaced reserved instances for now so I’d be tracking this issue https://github.com/hashicorp/terraform-provider-aws/issues/10785 if you are interesting in support for using Terraform to manage savings plans. post id: #48751593 label: provider
Adding sentences where the initial keywords (i.e. cost, bill, etc.) are mentioned provides additional insights:
Since licensing of SQL Server is expensive, I want to switch it off at least for the night. post id: #76474230 label: expense
I have multiple databases running in each environment which are charging me a lot of cost each month, so i wanted to downscale the DTUs to some lower count during non-working hours, again during working hours DTUs to be upscale back to actual DTUs count, it should happen automatically as per time settings every single day. post id: #74538381 label: cost
Some of these discussions point towards the need for architectural refactoring as a way of addressing cost-related concerns:
If you need zero cost during idle time, should you go with serverless with new design? post id: #56530721 label: cost
Aided by these highlighted sentences, we proceeded to manually inspect a representative sample of 37 questions [^15] https://www.calculator.net/sample-size-calculator.html?type=1&cl=95&ci=5&pp=2.6&ps=491&x=Calculate, considering that we selected 2.6% of a population of 19 1391913919\,13919 139 questions with Terraform tags. During the inspection, we aimed to verify whether: (1) the question indeed contained a cost-related discussion (either as the main problem or as the element of a description or argument), and (2) one of the properties is an essential element of the identified cost-related text. As a result, we found that all inspected questions were true positives. These combined observations show that the knowledge graph cannot only be used to identify actual pain points faced by cloud application developers, but it also points to the most recurrent ones.
Based on the co-occurrences identified in Figure 9, we proceeded to investigate if and how the knowledge graph could be augmented. We note that the frequency of the observed co-occurrences varies greatly while also being positively skewed, i.e., many low co-occurrences (min.: 1, Q1: 2, Q2: 5, Q3: 12, max.: 98). Following our goal of representing the most meaningful information in the knowledge graph, we processed the co-occurrence frequencies as follows:
  1. 1.
    mark preexisting graph edges that were not observed as absent;
  2. 2.
    classify the observed co-occurrences into present ($f\leq 2$), recurrent ($2<f\leq 4$), frequent ($4<f\leq 12$) and predominant ($12<f$) according to the four quartiles;
  3. 3.
    mark preexisting graph edges observed in the data according to their frequency class (established in the previous step); and
  4. 4.
    add new graph edges for predominant co-occurrences that were not present in the previous version.
The resulting knowledge graph is depicted in Figure 10. We note that the markings explained in the figure legend are aimed at helping the reader visualize the process. The augmented graph comprise all pre-existing edges with the addition of the new edges for predominant co-occurrences. Looking at the differences and co-occurrences frequencies, while many relationships between concepts are also observed in this dataset, we learn two main lessons: (a) the cost-related topics revolve mainly around the ability to try out different configurations (i.e., test) and updating configuration (i.e., change), also meaning that creating system alerts is not frequently discussed; and (b) discussions are more concrete in the sense that they get more specific about deployment matters such as features and benefits of providers, e.g., appropriate instances and policies. The latter point, in turn, led to a number of new relations being identified between “low level” concepts in the graph, e.g., between instance and policy or cluster.
@@ 6 Implications & Future Work
  • Use the extracted posts from Stack Overflow to extend the list of topics related to cost awareness, and update the knowledge graph of Figure 8 accordingly.
@@ 7 Threats to Validity
External Validity: Regarding external validity, the population in our data sets (from GitHub and Stack Overflow) may not represent all possible cost-related discussions in cloud projects. The inclusion of repositories hosted in other (closed-source) platforms, or that use other cloud orchestrators, or that use no cloud orchestrators, could lead to the identification of new discussions. The inclusion of discussion forums other than Stack Overflow could also lead to new discussions. However, our decisions were carefully considered with the aim of ensuring data quality, diversity and quantity while keeping the execution feasible for the available human resources. Furthermore, our goal was not to collect all possible evidence, to begin with, but rather to see if there is any evidence available. That said, the results triangulation (between GitHub- and Stack Overflow-based discussions) helps mitigate threats to external validity.
Construct Validity: The selection criteria, and most noticeably the defined keywords may threat the construct validity of our work. We mitigated this threat by piloting and testing our selection criteria, and considering the knowledge of both academic and industrial domain experts. Furthermore, the coding activity is naturally open to subjectivity and inconsistencies. To mitigate this threat, we followed a well-established process, and introduced an extra final step to consolidate the knowledge and systematically discuss the labels. Also, the topics derived from applying LDA may not fully represent their content. We mitigate this threat by manually inspecting the units of analysis during the investigation of RQ3, which was part of the axial coding and selective coding. This threat is also mitigated by the triangulation presented in Section 5. That said, we acknowledge that the Stack Overflow study itself suffers from its own threats to construct validity, mainly related to selection of appropriate data points, i.e., pertaining to cost-related discussion of Terraform-based deployments. The mitigation strategies in this case entail the use of Terraform-related tags that have been assigned by users, on top of using previously-validated keywords for searching cost-related entries within the filtered posts.